OpenStudy (anonymous):

Convert 600 L to grams, using dimensional analysis. PLEASE HELP! I am stuck on this whole conversion thing!

OpenStudy (anonymous):

Does it tell you what you have 600 Liters of? or maybe some conversion factor like density?

OpenStudy (anonymous):

The question is: You have a piece of copper that has a volume of 600 L. What is the mass of the sample, in units of grams?

OpenStudy (anonymous):

ok give me a sec to find the desnity of copper and i will set it up for you

OpenStudy (anonymous):

It is trying to set it up, on the page, using dimensional analysis. It has a drag and drop box, to put the right parts where they go and I cannot seem to figure it out. This is a very HARD subject for me unfortunately.

OpenStudy (anonymous):

\[600 L Cu (\frac{1000ml Cu}{1 L Cu})(\frac{8.96 g Cu}{1 ml Cu})=5.376x10^6 g Cu\] this is based off the density being 8.96 g/ml we must remember that 1ml=1cm^3